OpenStudy (anonymous):

F=kmn/r^2 How do I solve for M

OpenStudy (anonymous):

F = kmn/r^2 multiply both side by r^2 so you can get rid of the r^2 on the bottom r^2F = kmn divide by kn to get M by itself r^2F ---- = m kn That is if the r^2 is under the entire kmn kmn ---- r^2

OpenStudy (anonymous):

that is correct it is under the entire kmn so the answer is r^2F/kn

OpenStudy (anonymous):

so if it is F=-kmn/r^2 my answer would be - R^2F/kn

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay that was easier than I thought. Thank you so much for the help

OpenStudy (anonymous):

Rouault are you still here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{3y+4=4}\] the solution is y= ?